Sipho Jackson and his wife Joy Jackson entered a competition to win a Mercedes A-Class worth R700 000 from their nearest dealer. Sipho sent in 40 entries while Joy sent in 10 entries. The dealership administrators revealed that only 200 entries qualified. Use this information to answer the following questions:

1.1 What is the probability that Sipho will win first prize? (2)

1.2 If Sipho wins the first prize, what is the probability that Joy will win the second prize? (Assuming that the first prizewinning entry is not put back into the dealer’s container with the entry forms). (2)

1.3 What is the probability that Sipho and Joy will win first and second prizes? (3)


1
Expert's answer
2021-07-30T08:51:24-0400

Question 1.1

P (Sipho) = 40/200

              = 1/5

Question 1.2

1-1/5 = 4/5

Question 1.2

1/5 X 4/5=4/25